Find the transitive closure A={1,2,3,4,5}

R={ (2,3),(3,5),((1,2),(2,5),(1,3),(4,5),(5,2)} by using Warshall’s Algorithm.


1
Expert's answer
2021-09-07T19:06:18-0400

M(R)="\begin{pmatrix}
 0 & 1 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 & 1 \\ 
 0 & 0 & 0 & 0 & 1 \\
 0 & 0 & 0 & 0 & 1 \\
 0 & 1 & 0 & 0 & 0 \\
\end{pmatrix}" matrix of relation R

Step 1: k=1

previous: "\begin{pmatrix}
 \bar 0 & \bar 1 & \bar 1 & \bar0 & \bar0 \\
\bar 0 & 0 & 1 & 0 & 1 \\
 \bar0 & 0 & 0 & 0 & 1 \\
 \bar0 & 0 & 0 & 0 & 1 \\
 \bar0 & 1 & 0 & 0 & 0 \\
\end{pmatrix}" next : "\begin{pmatrix}
 0 & 1 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 & 1 \\ 
 0 & 0 & 0 & 0 & 1 \\
 0 & 0 & 0 & 0 & 1 \\
 0 & 1 & 0 & 0 & 0 \\
\end{pmatrix}"

Matrix unchanged

Step 2: k=2

previous:"\begin{pmatrix}
 0 & \bar1 & 1 & 0 &> 0 \\
 \bar0 &\bar 0 & \bar1 & \bar0 & \bar1 \\ 
 0 & \bar0 & 0 & 0 & 1 \\
 0 & \bar0 & 0 & 0 & 1 \\
 0 & \bar1 & >0 & 0& >0 \\
\end{pmatrix}" next: "\begin{pmatrix}
 0 & 1 & 1 & 0 & 1 \\
 0 & 0 & 1 & 0 & 1 \\ 
 0 & 0 & 0 & 0 & 1 \\
 0 & 0 & 0 & 0 & 1 \\
 0 & 1 & 1 & 0& 1 \\
\end{pmatrix}"

Thee edges have added.

Step 3: k=3

previous:"\begin{pmatrix}
 0 & 1 & \bar 1 & 0 & 1 \\
 0 & 0 & \bar 1 & 0 & 1 \\ 
 \bar0 & \bar 0 & \bar0 & \bar0 & \bar 1\\
 0 & 0 & \bar0 & 0 & 1 \\
 0 & 1 & \bar1 & 0& 1 \\
\end{pmatrix}" next: "\begin{pmatrix}
 0 & 1 & 1 & 0 & 1 \\
 0 & 0 & 1 & 0 & 1 \\ 
 0 & 0 & 0 & 0 & 1\\
 0 & 0 & 0 & 0 & 1 \\
 0 & 1 & 1 & 0& 1 \\
\end{pmatrix}"

Matrix is unchanged

Step 4: k=4 Matrix will not change because fourth column is zeros.

Step 5: k=5

"previous:\begin{pmatrix}

 0 & 1 & 1 & 0 & \bar1 \\

 0 & >0 & 1 & 0 & \bar1 \\ 

 0 & > 0 & >0 & 0 & \bar1\\

 0 & >0 &> 0 & 0 & \bar1 \\

 \bar0 & \bar1 & \bar1 &\bar 0& \bar1 \\

\end{pmatrix}"

final: "\begin{pmatrix}

 0 & 1 & 1 & 0 & 1 \\

 0 & 1& 1 & 0 & 1 \\ 

 0 & 1 & 1& 0 & 1\\

 0 & 1 &1 & 0 & 1 \\

 0 & 1 & 1 & 0& 1 \\

\end{pmatrix}" - matrix of transitive closure

Thus "\bar R" ={(1,2),(1,3),(1,5),(2,2),(2,3),(2,5),(3,2),(3,3),(3,5),(4,2),(4,3),(4,5),(5,2),(5,3),(5,5)}